# The First Photometric Evidence of a Transient/Variable Source at 5 with JWST

### v1

@@ 4.3 Tidal Disruption Event
A TDE occurs when a supermassive black hole (SMBH) shreds apart a star that passes within its tidal radius, generating a luminous accretion flare (Rees, 1988; Evans & Kochanek, 1989; Ulmer, 1999). TDEs are expected to be abundant at high redshifts, due to the high central stellar densities of high redshift galaxies and the slightly lower SMBH masses (Karmen et al., 2025). Inayoshi et al. (2024) find that deep JWST surveys can discover TDEs occurring in moderately obscured AGN (e.g., with typical attenuation levels of mag, as seen in LRDs) up to 4–7 even with small survey areas. These observations enable the exploration for the mass distribution of SMBHs at the low-mass end, where their abundance is sufficiently high for detection. Pfister et al. (2021) uses hydrodynamical simulations to show that the TDE rate can be enhanced at early cosmological times, and that TDEs could be a viable growth mechanism for massive black hole seeds. Additionally, Karmen et al. (2025) reports the discovery of a high-redshift TDE candidate in the COSMOS-Web field, identified using novel color and morphology selection criteria. This further suggests enhanced TDE rates at high-redshift relative to the local rate.
@@ 5 Conclusion
We discovered AT 2023adya by differencing F356W NIRCam imaging in GOODS-N taken one observer-frame year apart by the JADES and CONGRESS programs. AT 2023adya is thus far the highest-redshift transient/variable source showing a change in photometric brightness discovered with JWST, residing in a 5.274 galaxy which faded from 26.050.02 mag to 26.240.02 mag (+0.19 mag) over approximately two rest-frame months. The F356W difference image shows a clear residual signal ( 28.010.17 mag). AT 2023adya does not, however, exhibit rest-frame UV variability in the F090W or F115W filters. While we cannot determine the true nature of AT 2023adya, we explore various sources that may create the observed brightness change, including an SN explosion, merger-induced AGN activity, and a TDE.
@@
AT 2023adya’s host is a member of a high-confidence galaxy merger pair (Puskás et al., 2025), indicating that AT 2023adya may be the signature of merger-induced AGN activity. However, we did not observe broad H emission in AT 2023adya’s F444W NIRCam/grism spectrum one observer-frame week after the Epoch1 imaging, meaning that AT 2023adya is unlikely to be a Type 1 AGN. This disfavors the AGN scenario, as Type 2 AGN exhibit negligible variability. NIRCam/Grism’s limited depth does make it difficult to exclude the existence of a faint broad-line component in the F444W spectrum, but a 19% brightness change would have made such a component visible in the spectrum.
@@
Although we cannot determine AT 2023adya’s true nature with the available data, this discovery is significant because it represents the first piece of photometric evidence that JWST can robustly detect transient/variable sources at 5. It is notable that the JTS discovered no such sources at 5 in GOODS-S. Had a source similar to AT 2023adya been in the JADES Deep Field within GOODS-S, the JTS selection criteria likely would have identified it. The lack of 5 transient and variable sources in the JTS sample indicates the rarity of discovering such sources at high redshift when the baseline between observations is only one observer-frame year. To continue building the 5 transient and variable source sample, which may include exotic sources such as Population III SNe, we will conduct multi-epoch multi-filter JWST observations with baselines longer than one observer-frame year through Cycle-4 JWST Program 8060 (PI: Egami).
@@ Acknowledgements
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1181, 1895 and 3577. The specific JWST observations analyzed can be accessed via https://dx.doi.org/10.17909/8tdj-8n28 (catalog doi:10.17909/8tdj-8n28) and http://dx.doi.org/10.17909/6rfk-6s81 (catalog doi:10.17909/6rfk-6s81 ). Support for program #3577 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. The authors acknowledge the FRESCO team for developing their observing program with a zero-exclusive-access period.

### v2

@@ 4.3 Tidal Disruption Event
A TDE occurs when a supermassive black hole (SMBH) shreds apart a star that passes within its tidal radius, generating a luminous accretion flare (Rees, 1988; Evans & Kochanek, 1989; Ulmer, 1999). TDEs are expected to be abundant at high redshifts due to the high central stellar densities of high redshift galaxies and the slightly lower SMBH masses (Karmen et al., 2025). Inayoshi et al. (2024) find that deep JWST surveys can discover TDEs occurring in moderately obscured AGN (e.g., with typical attenuation levels of mag, as seen in LRDs) up to 4–7 even with small survey areas. These observations enable the exploration of the mass distribution of SMBHs at the low-mass end, where their abundance is sufficiently high for detection. Pfister et al. (2021) uses hydrodynamical simulations to show that the TDE rate can be enhanced at early cosmological times, and that TDEs could be a viable growth mechanism for massive black hole seeds. Additionally, Karmen et al. (2025) reports the discovery of a high-redshift TDE candidate in the COSMOS-Web field, identified using novel color and morphology selection criteria. This further suggests enhanced TDE rates at high-redshift relative to the local rate.
@@ 5 Conclusion
We discovered AT 2023adya by differencing F356W NIRCam imaging in GOODS-N taken one observer-frame year apart by the JADES and CONGRESS programs. AT 2023adya is thus far the highest-redshift transient/variable source showing a change in photometric brightness discovered with JWST, residing in a 5.274 galaxy which faded from 26.050.02 mag to 26.240.02 mag (+0.19 mag) over approximately two rest-frame months. The F356W difference image shows a clear residual signal ( 28.010.17 mag). AT 2023adya does not, however, exhibit a rest-frame UV brightness change in the F090W or F115W filters. While we cannot determine the true nature of AT 2023adya, we explore various sources that may create the observed brightness change, including an SN explosion, AGN, and a TDE.
@@
We cannot rule out the possibility that AT 2023adya is a Type 1 AGN whose continuum is dominated by star formation. In this case, it is likely that the broad H component would be too faint for detection in the F444W NIRCam/Grism spectrum. Additionally, we cannot rule out the possibility that AT 2023adya is a “true” or “naked” Type 2 AGN, but we do not favor this possibility due to the rarity of these objects.
@@
Although we cannot determine AT 2023adya’s true nature with the available data, this discovery is significant because it represents the first piece of photometric evidence that JWST can robustly detect transient/variable sources at 5. It is notable that the JTS discovered no such sources at 5 in GOODS-S. Had a source similar to AT 2023adya been in the JADES Deep Field within GOODS-S, the JTS selection criteria likely would have identified it. The lack of 5 transient and variable sources in the JTS sample indicates the rarity of discovering such sources at high redshift when the temporal baseline between observations is only one observer-frame year. To continue building the 5 transient and variable source sample, which may include exotic sources such as Population III SNe, we will conduct multi-epoch multi-filter JWST observations with baselines longer than one observer-frame year through Cycle-4 JWST Program 8060 (PI: Egami).
@@ Acknowledgements
We thank the anonymous referee whose comments and suggestions led to significant improvements in the paper. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1181, 1895 and 3577. The HLF, FRESCO, JADES, and CONGRESS data used in this paper can be found on MAST (HLF Team, 2015; FRESCO Team, 2023; JADES Team, 2024; CONGRESS Team, 2025). Support for program #3577 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. The authors acknowledge the FRESCO team for developing their observing program with a zero-exclusive-access period.